OpenStudy (anonymous):

Someone please help me, i have no clue what to do!!! Perform the indicated operations.

OpenStudy (anonymous):

\[\frac{ x+1 }{ 3y }+\frac{ x-2 }{ 4y }-\frac{ x+3 }{ 6y }\]

OpenStudy (anonymous):

get a common denominator

OpenStudy (anonymous):

ok to 12y?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

then just multiply top and bottom of each fraction by the number that makes the bottom 12y, then combine like terms on top

OpenStudy (anonymous):

so (4x+4)+(3x-6)-(2x+6) then 5x+4/12y

OpenStudy (anonymous):

yo got it!

OpenStudy (anonymous):

Awesome, thank you veery much, i couldn't figure out how to start it out.

OpenStudy (anonymous):

and that is in simplest form since the variables are different on top and bottom

OpenStudy (anonymous):

yw! np, starting out is always the hardest step

OpenStudy (anonymous):

Yeah, it always is